# Laser mode-hopping assisted all-optical single beam pulsed atomic magnetometer

### v1

@@ Abstract
We demonstrate an all-optical single beam pulsed atomic magnetometer assisted by laser mode-hopping in a distributed Bragg reflector (DBR) laser. We implement a temporal sequence of the laser current; sinusoidal current modulation including the laser mode-hop current for synchronous optical pumping and the following constant current for paramagnetic Faraday rotation measurements to probe the free induction decay (FID) of transverse 87Rb spin polarization. Repetitive sudden frequency shifts of 20 GHz around the pressure-broadened 87Rb spectra originating from laser mode-hopping enables discontinuous optical pumping modulation with a large depth which enhances transverse spin polarization. We achieve a sensitivity of 3.77 pT/Hz1/2 in a magnetic field of 14 $\mu$T, limited by the performance of the frequency counter. The Cramer-Rao lower bound (CRLB) of the sensitivity due to the non-magnetic noise such as photon shot-noise is 191 fT/Hz1/2. Our approach based on laser mode-hopping can be applied to miniaturization of all-optical atomic magnetometers with sub-pT/Hz1/2 sensitivities.
@@ I Introduction
Laser mode-hopping refers to a jump of the lasing mode, which stems from the relative frequency shift of longitudinal cavity modes to the gain curve originating from temperature or injection current changes. It is generally considered to be problematic in laser spectroscopy and telecommunication because it accompanies unwanted intensity fluctuations and large sudden frequency shifts. Many researches have been performed to extend the mode-hop free range to facilitate continuous wavelength tuning gong2014mode ; dutta2012mode ; fuhrer2008extension . In the aspect of understading underlying physics of mode-hopping, there also have been attempts to study dynamics of laser mode-hopping kundu2018ultrafast ; butler2017direct ; happach2020effect .
@@
A large sudden frequency shift feature during laser mode-hopping, however, may be useful in some applications; in Bell-Bloom type atomic magnetometer experiments where the modulation of optical pumping rate to induce spin-precession is involved, large sudden frequency shifts during mode-hopping offers enhancement of spin-polarization by increasing the modulation depth of optical pumping. In order to build up large transverse spin polarization in all-optical Bell-Bloom type atomic magnetometers, a variety of techniques including sinusoidal modulation of the laser frequency has been explored kimball2017situ ; jimenez2009sensitivity . Some advancements in nonlinear magneto-optical rotation (NMOR) magnetometer were presented based on pulse-shaped amplitude modulation of beam, taking advantage of the discontinuous modulation gawlik2006nonlinear ; gerginov2017pulsed . Obtaining a discontinuous lineshape in the modulation function, however, requires cumbersome optics like acousto-optic modulators (AOM) or intricate techniques to generate square-wave RF magnetic field ranjbaran2017effects ; hunter2018free .
@@
Frequency switching by means of mode-hopping, on the other hand, only requires weak sinusoidal modulation of the laser diode (LD) injection current to create a large discontinuous optical pumping modulation. Such an effectiveness can be estimated from the optical pumping rate formula. The optical pumping rate is a function of laser intensity and detuning, given by $R\propto I/(\Delta\nu^{2}+(\Gamma_{L}/2)^{2})$ where $I$ is the laser intensity, $\Delta\nu$ is the laser detuning, and $\Gamma_{L}$ is the line broadening. For example, when the frequency shift during a mode-hopping is $2\Gamma_{L}$, the optical pumping rate of $\Delta\nu=2\Gamma_{L}$ is 17 times smaller than that of $\Delta\nu=0$. To acheive the same modulation depth by using an intensity modulation, the laser intensity have to be changed from $I$ to $I/17$. To obtain such a large modulation depth an external light modulator, for example, AOM or electro-optic modulator (EOM), can be applied. For mininarization of the device, direct injection current modulation into the LD can be adopted instead of using an external light modulator. Considering the linear frequency shift of the LD as a function of injection current, in our case, -1 GHz/mA in the mode-hop free range, an amplitude of current modulation of 20 mA is required to reach a 20 GHz detuning corresponding to a double of $\Gamma_{L}=10$ GHz, a typical line broadening of the Rb D1 line in nitrogen buffer gas. Such a detuning can be easily obtained by only a few mA when laser mode-hopping is adopted. Moreover, the injection current into the LD is not needed to be discontinuous. A continuous laser modulation with a small modulation depth leads to a large depth of discontinuous optical pumping modulation when laser mode-hopping is applied. It is an advantage of laser mode-hopping assisted optical pumping because a large amplitude of discontinuous laser injection current can be a source of the unwanted magnetic field noise.
@@
Frequency switching feature can be also helpful for implementing single beam all-optical atomic magnetometers based on FID measurements. We can easily shift the laser frequency to be far detuned from or to be resonant on the atomic resonance by utilizing laser mode-hopping. A temporal sequence of the laser current containing the mode-hop current enables FID measurements. In the on-resonance period where the laser is resonant on the atomic transition, the atomic spin is polarized along the laser beam direction. In the following off-resonance period where the laser is far detuned from the atomic transition due to the mode-hopping, optical pumping ceases and the FID of the transversal atomic spin is induced. The far-detuned laser undergoes the paramagnetic Faraday rotation depending on the transversal atomic spin which precesses with the Larmor frequency.
@@
Usually, single beam atomic magnetometers utilize a magnetic field modulation technique using a coil liu2020single ; ding2020single ; oelsner2022integrated . Their structure are very simple but they can suffer from magnetic field interference when forming a gradiometer or a magnetometer array. To diminish such the effect of magnetic interference all-optical atomic magnetometers have been developed widely. All-optical vector atomic magnetometer was reported patton2014all . An all-optical atomic magnetometer with a co-linear configuration to measure the biomedical magnetic field was implemented by placing the pump and probe lasers at a slight angle within the same window limes2020portable . Single beam all-optical atomic magnetometers based on an external light modulator were studied petrenko2022towards ; hunter2018free . Even, all-optical single beam atomic magnetometer based on direct current modulation to the vertical-cavity surface-emitting laser (VCSEL) was already demonstrated with a sensitivity level of 10 pT/Hz1/2 jimenez2009sensitivity ; hunter2018free . However, the low power of the VCSEL is insufficient to obtain a significant optical pumping rate $R(t)$ so that the sensitivity is limited. Our strategy for single beam all-optical atomic magnetometer without the use of an external light modulator is that adopting a DBR laser instead of a VCSEL to acheive higher optical pumping rate and applying current sequence including a laser mode-hopping, which consists of the optical pumping stage and the following probing stage to measure the FID of macroscopic 87Rb spin.
@@
In this paper, we demonstrate an all-optical single beam pulsed atomic magnetometer utilizing a mode-hopping of a DBR laser. We employ mode-hopping to selectively switch the laser frequency between on-resonance and off-resonance with the 87Rb D1 line so that a sequence of optical pumping and the FID probing is realized. We observe the FID as a function of offset, modulation amplitude of current modulation and analyze the discontinuous optical pumping effect. Additionally, we measure the sensitivity of our magnetometer and compare it to the CRLB due to the non-magnetic noise which shows the reachable sensitivity of our magnetometer.
@@ II Model for discontinuous optical pumping
Figure 1: Illustration of the concept of frequency modulation assisted by laser mode-hopping. A sinusoidal modulation of laser injection current is converted to a discontinuous output frequency profile. The red and blue area represent the adjacent cavity modes in a DBR laser. A small variation around the mode-hop current results in a huge shift on laser output frequency. The frequencies of these modes correspond to the resonant and off-resonant condition of pressure broadened 87Rb absorption profile, respectively. The repetitive jump between cavity modes induces a discontinuous synchronous optical pumping.
A discontinuous optical pumping profile is achievable by incorporating a simple sinusoidal modulation of the laser injection current with assistance of a mode-hopping (see Figure 1). When the sinusoidal current modulation contains the mode-hop current, the laser frequency jumps by the frequency difference between cavity modes. Assumming that the frequency of the mode 1 (or the mode 2) is resonant on (or far-detuned from) the atomic resonance, the optical pumping rate, $R\propto I/(\Delta\nu^{2}+(\Gamma_{L}/2)^{2})$, shows steep changes as shown in Fig. 1. The change on optical pumping rate can be regarded to be sufficiently sudden, in other words, discontinous because the swithcing time of mode-hopping given by an order of 10 ns is much shorter than the Larmor period of 2.85 $\mu$s in a magnetic field of $50~{}\mu$T butler2017direct .
In order to analyze the effect of the discontinuous optical pumping induced by a mode-hopping, we investigate the macroscopic spin Bloch equation including the optical pumping modulation term, $R(t)$. The Bloch equation model, widely used for atomic magnetometers, is suitable for modeling arbitrary modulations distorted by mode-hopping grujic2013atomic . The Bloch equation is given by
@@
Here, $\vec{s}=(e^{*}\times e)/i$ represents the photon spin vector where $e$ represents a unit polarization vector in the yz plane so that $\vec{s}=s\hat{x}$. $\Gamma_{re}$ is the spin relaxation rate mainly contributed by the spin exchange collision between 87Rb atoms. $R(t)$ is the optical pumping rate for unpolarized atoms, depending on the laser detuning from the $D_{1}$ resonance and the laser intensity vasilakis2011precision . $\gamma$ is the gyromagnetic ratio. We consider the magnetic field is directed in the $\hat{y}$ axis.
@@
In the case of square wave optical pumping modulation, $R_{0}$ varies accordingly to the duty cycle $d$ when the maximum value of $R(t)$ is fixed. Assuming the maximum optical pumping rate of the square wave optical pumping modulation is $R_{m}$, then $S_{0}$ can be expressed by
@@
The maximum $S_{0}$ is 1/2 when $d\rightarrow 0$ in the regime that $\Gamma_{re}\ll R_{m}d$. This implies that a delta-function like optical pumping induces the largest $S_{0}$ as long as the peak intensity of a laser pulse is strong enough to satisfy $R_{m}\gg\Gamma_{re}/d$. The reason why the delta-function like optical pumping is optimal can be described in terms of $\vec{s}$ and $\langle\vec{S}\rangle$. In the Bell-Bloom configuration, optical pumping not only creates the Rb spin polarization along the direction of the pumping beam but also relaxes the transverse spin. When the direction of $\vec{s}$ does not coincide with $\langle\vec{S}\rangle$, optical puming destructs the transverse spin precession and aligns the atomic spin with the direction of the pumping beam. It ceases to serve as a transverse spin relaxation mechanism only if $\langle\vec{S}\rangle$ is oriented in the same direction of $\vec{s}$. As a result, we can conclude that the delta-function like optical pumping is optimal. In the regime that the optical pumping rate $R_{m}$ is not much larger than $\Gamma_{re}/d$, the nonzero $d$ can be optimal. For example, the weak pumping condition that $\Gamma_{re}=2\pi\times 685$ Hz, $R_{m}=2\pi\times 300$ Hz, and $|s|=1/2$, the optimal duty cycle is caculated to be around 50 %. In this case, $S_{0}$ of the discontinuous optical pumping of a 50 % duty cycle is $4/\pi$ times larger than that of the sinusoidal optical pumping modulation if the maximum optical pumping rates are same.
@@ III Experimental Setup
Figure 2: Experimental setup for the single beam all-optical magnetometer. DBR laser: Distributed Bragg Reflector laser, PBS: Polarization Beam Splitter, QWP: Quarter Wave Plate, HWP: Half Wave Plate, BPD: Balanced Photodetector. A burst sine wave modulation to the DBR laser relizes pulsed operation in 1 kHz. Red and blue regions show pumping and probing stage of pulsed magnetometer, respectively.
@@
Figure 3: (a) 87Rb D1 absorption spectrum in the atomic vapor cell. The grey dots indicate the absorption of the linearly polarized laser beam. The black solid line depicts the fitting curve. The Full Width at Half Maximum linewidth $\Gamma_{L}$ is 10.5 GHz which implies that the pressure of $\textrm{N}_{2}$ in our cell is about 580 Torr. The red and blue solid lines depict the absorption spectrum of hyperfine transitions from $F_{g}=1$ and $F_{g}=2$, respectively. The black arrows show how the laser frequency detuning changes as the laser current grows. (b) Laser detuning as a function of LD injection currents. The mode-hop current is 67.1 mA. (c) Relative Faraday rotation angle (black solid line) and transverse spin relaxation rate (red solid line) $\Gamma_{re}$ as a function of LD injection current. The injection current below the mode-hop current induces a larger Faraday rotation angle and transverse spin relaxation rate because the corresponding laser frequency is within the 87Rb D1 absorption spectrum. (d) Relative optical power of the laser beam after passing through the atomic vapor cell as a function of LD injection current. The laser injection current above the mode-hop current gives higher beam power because the off-resonant light expriences weak absorption.
@@
The experimental setup is shown in Figure 2. A DBR laser (Photodigm 794.978 nm series with TOSA package) is used as a light source. Mode-hopping in the DBR laser occurs under fixed condition due to it’s monolithic nature, allowing the jump to the output frequency controllable. Two main factors influence the frequency tuning of the DBR laser output: the temperature of the LD heat sink and the laser injection current. When modulating injection current with a stabilized laser temperature, a slight alteration in laser current triggers a transition between two adjacent cavity modes as illustrated in Fig. 1. The transition of lasing mode near the mode-hop current results in a significant jump in the laser output frequency, which was almost 20 GHz in our case. A sinusoidal modulation of LD injection current with aid of a mode-hopping, therefore, is converted to a discontinuous laser output frequency profile without any extra optics.
@@
The linearly polarized laser beam is first transformed into an elliptically polarized beam after passing through a QWP. The optimal QWP angle to increase the signal is 25 and the correspoding $|s|$ is about 0.67 (see the appendix). The laser beam serves both as the probing and pumping beam. It illuminates a $5\times 5\times 5$ m3 cubic cell which contains 87Rb vapor and $\mathrm{N_{2}}$ gas at a pressure of 580 Torr (see Figure 3 (a)).
@@
The home-made atomic gas cell is heated up to 103 C using a dual-layer structured pair of AC heaters designed to negate magnetic field noise yim2018note . After passing through the atomic gas cell, which is inside a two layered $\mu$-metal shield, the laser beam goes through a HWP. A polarimeter, consisting of a PBS and a BPD, measures the degree of polarization rotation of elliptically polarized beam induced by the paramagnetic Faraday rotation effect. The rotation angle of the polarization is proporitonal to $\Delta\nu/[\Delta\nu^{2}+(\Gamma_{L}/2)^{2}]$ where $\Delta\nu$ is the laser detuning in the probing stage.
@@
Our pulsed magnetometer alternates pumping and probing mode at a repetition rate of 1 kHz. The laser injection current is modulated by a burst sine wave generated by a function generator (Siglent’s SDG2082X), realizing a pulsed operation of magnetometer. In the pumping stage, depicted as red region in Fig. 2, the LD injection current is modulated at the Larmor frequency of ${}^{87}Rb$ (which is 100 kHz in our case), incorporating a mode-hop current into the modulation of the LD injection current. The mode-hop current is 67.1 mA when the TEC temperature of the LD is set to 32.05 C. When the LD injection current is above the mode-hop current, the laser output frequency deviates from the Rb resonance line as laser detuning becomes almost 20 GHz, shown as an arrow in Fig. 3. On the other hand, when the LD current is below the mode-hop current, the laser output frequency gets close to the $F_{g}=1$ transition line as laser detuning becomes almost 0 GHz. Due to the pressure broadening of 10.5 GHz, the laser beam also weakly induces the transitions from $F_{g}=2$.
@@
In the probing stage, illustrated by the blue region in Fig. 2, the sinusoidal modulation of the LD injection current stops and injection current is fixed to a certain value at which the laser frequency is off-resonant. The detuned laser beam undergoes the paramagnetic Faraday rotation effect where the polarization rotation angle is proportional to the atomic spin projected in the direction of the laser beam propagation. The laser beam with rotated polarization is directed into a polarimeter. A damped sine signal on which the FID is imprinted can be obtained from the BPD. It can be described by
@@
In order to estimate the sensitivity of our magnetometer, we apply a frequency counter (53230A, Keysight) which measures the frequency of the signal based on the zero-crossing detection method. The reciprocal of time between two zero-crossing points becomes the Larmor frequency.
Figure 4: (a) FID amplitude as a function of LD current DC offsets and modulation depths. The results are separated into four regions divied by two dashed line (the boundaries where modulation includes mode-hop current) and straigt line(the mode-hop current at 67.1 mA). (b) Semi-theoretical result. $S_{0}$ is calculated based on the optical pumping profile shown in the top pannel. The optical pumping rate profile is extracted by subtracting minimum relaxation rate from the measured relaxation rate shown at Fig. 3 (c). (c) Semi-theoretical result based on the step-function like optical pumping profile. (d) FID signals at four regions. The insets show 1-period temporal pumping rate profiles extracted from the top pannel of (b). The modulation depth is fixed to 1.8 mApp.

### v2

@@ Abstract
We demonstrate an all-optical single beam pulsed atomic magnetometer assisted by laser mode-hopping in a distributed Bragg reflector (DBR) laser. We implement a temporal sequence of the laser current, with sinusoidal current modulation including the laser mode-hop current for synchronous optical pumping, and a following constant current for paramagnetic Faraday rotation measurements, to probe the free induction decay (FID) of transverse 87Rb spin polarization. Repetitive sudden frequency shifts of 20 GHz around the pressure-broadened 87Rb spectra, originating from laser mode-hopping, enable discontinuous optical pumping modulation with a large depth, which enhances transverse spin polarization. We achieved a sensitivity of 0.6 pT/Hz1/2 in a magnetic field of 27 $\mu$T, mainly limited by the photon-shot-noise and the magnetic field noise induced by the current noise in the current supply for driving the bias magnetic field coil. The Cramer-Rao lower bound (CRLB) of the sensitivity due to the non-magnetic noise such as photon shot-noise is 131 fT/Hz1/2. Our approach based on laser mode-hopping can be applied for the miniaturization of all-optical atomic magnetometers with sub-pT/Hz1/2 sensitivities.
@@ I Introduction
Laser mode-hopping refers to a jump in lasing mode which stems from a relative frequency shift in longitudinal cavity modes to the gain curve. It originates from changes in temperature or injection current and is generally considered to be problematic in laser spectroscopy and telecommunications because it is accompanied by unwanted intensity fluctuations and large sudden frequency shifts. Many studies have attempted to extend the mode-hop free range to facilitate continuous wavelength tuning gong2014mode ; dutta2012mode ; fuhrer2008extension . There have also been attempts to study the dynamics of laser mode-hopping to better understand the underlying physics of mode-hopping. kundu2018ultrafast ; butler2017direct ; happach2020effect .
@@
A large sudden frequency shift feature during laser mode-hopping, however, may be useful in some applications. For example, in Bell-Bloom type atomic magnetometer experiments where the modulation of optical pumping rate to induce spin-precession is involved, large sudden frequency shifts during mode-hopping can enhance spin-polarization by increasing the modulation depth of the optical pumping. In order to build up large transverse spin polarization in all-optical Bell-Bloom type atomic magnetometers, a variety of techniques have been explored, including sinusoidal modulation of the laser frequency kimball2017situ ; jimenez2009sensitivity . Some advancements in the development of the nonlinear magneto-optical rotation (NMOR) magnetometer have been reported, based on pulse-shaped amplitude modulation of the beam, which takes advantage of the discontinuous modulation gawlik2006nonlinear ; gerginov2017pulsed . Obtaining a discontinuous lineshape in the modulation function, however, requires cumbersome optics such as acousto-optic modulators (AOM) or intricate techniques to generate a square-wave RF magnetic field ranjbaran2017effects ; hunter2018free .
@@
Frequency switching by means of mode-hopping, on the other hand, only requires weak sinusoidal modulation of the laser diode (LD) injection current to create a large discontinuous optical pumping modulation. Its effectiveness can be estimated from the optical pumping rate formula. The optical pumping rate is a function of laser intensity and detuning, given by $R\propto I/(\Delta\nu^{2}+(\Gamma_{L}/2)^{2})$ where $I$ is the laser intensity, $\Delta\nu$ is the laser detuning, and $\Gamma_{L}$ is the line broadening. For example, when the frequency shift during mode-hopping is $2\Gamma_{L}$, the optical pumping rate of $\Delta\nu=2\Gamma_{L}$ is 17 times smaller than that of $\Delta\nu=0$. To achieve the same modulation depth using intensity modulation, the laser intensity has to be changed from $I$ to $I/17$. To obtain such a large modulation depth, an external light modulator, for example, AOM or electro-optic modulator (EOM), can be applied. To miniaturize the device, current modulation can be directly injected into the LD instead of using an external light modulator. Since the linear frequency shift of the LD is a function of injection current, in our case, -1 GHz/mA in the mode-hop free range, an amplitude of current modulation of 20 mA is required to achieve a 20 GHz detuning. This corresponds to a doubling of $\Gamma_{L}=10$ GHz, which is a typical line broadening of the Rb D1 line in nitrogen buffer gas. Such detuning can be easily obtained with only a few mA when laser mode-hopping is adopted. In addition, the injection current into the LD does not need to be discontinuous. A continuous laser modulation with a small modulation depth leads to a large depth of discontinuous optical pumping modulation when laser mode-hopping is applied. This is an advantage of laser mode-hopping assisted optical pumping, because a large amplitude of discontinuous laser injection current can be a source of unwanted magnetic field noise.
@@
The frequency switching feature can also be helpful for implementing single beam all-optical atomic magnetometer based on FID measurements. By utilizing laser mode-hopping we can easily shift the laser frequency to be far detuned from, or to be resonant with, the atomic resonance. A temporal sequence of the laser current containing the mode-hop current enables FID measurements. In the on-resonance period when the laser is resonant with the atomic transition, the atomic spin polarizes along the direction of the laser beam. During the following off-resonance period, when the laser is far detuned from the atomic transition due to the mode-hopping, optical pumping stops and induces the FID of the transverse atomic spin. The far-detuned laser experiences paramagnetic Faraday rotation depending on the transverse atomic spin that precesses with the Larmor frequency.
@@
Usually, single beam atomic magnetometers utilize a magnetic field modulation technique using a coil liu2020single ; ding2020single ; oelsner2022integrated . Their structure is very simple but they can suffer from magnetic field interference when forming a gradiometer or a magnetometer array. To diminish the effect of magnetic interference, all-optical atomic magnetometers have been widely developed. An all-optical vector atomic magnetometer was reported patton2014all . An all-optical atomic magnetometer with a co-linear configuration to measure biomedical magnetic fields was implemented by placing the pump and probe lasers at a slight angle within the same window limes2020portable . Single beam all-optical atomic magnetometers based on an external light modulator have also been studied petrenko2022towards ; hunter2018free . And an all-optical single beam atomic magnetometer based on direct current modulation to the vertical-cavity surface-emitting laser (VCSEL) has already been demonstrated with a sensitivity level of 10 pT/Hz1/2 jimenez2009sensitivity ; hunter2018free . However, the low power of the VCSEL is insufficient and unable to achieve a significant optical pumping rate $R(t)$ so that its sensitivity is limited.
Our strategy for a single beam all-optical atomic magnetometer without the use of an external light modulator involves adopting a DBR laser instead of a VCSEL to achieve a higher optical pumping rate, and applying a current sequence including laser mode-hopping. It consists of an optical pumping stage and a following probing stage to measure the FID of macroscopic 87Rb spin.
@@
In this paper, we demonstrate an all-optical single beam pulsed atomic magnetometer utilizing the mode-hopping of a DBR laser. We employ mode-hopping to selectively switch the laser frequency between on-resonance and off-resonance with the 87Rb D1 line so that a sequence of optical pumping and FID probing is realized. We observe the FID as a function of offset, modulation amplitude of current modulation and analyze the discontinuous optical pumping effect. Additionally, we measure the sensitivity of our magnetometer and compare it to the CRLB considering non-magnetic noise, demonstrating the achievable sensitivity of our magnetometer.
@@ II Model for discontinuous optical pumping
Figure 1: Illustration of the concept of frequency modulation assisted by laser mode-hopping. A sinusoidal modulation of laser injection current is converted to a discontinuous output frequency profile. The red and blue areas represent the adjacent cavity modes in a DBR laser. A small variation around the mode-hop current results in a huge shift in laser output frequency. The frequencies of these modes correspond to the resonant and off-resonant conditions of the pressure-broadened 87Rb absorption profile, respectively. The repetitive jumps between cavity modes induces discontinuous synchronous optical pumping.
A discontinuous optical pumping profile can be achieved by incorporating a simple sinusoidal modulation of the laser injection current with assistance of mode-hopping (see Figure 1). When the sinusoidal current modulation contains the mode-hop current, the laser frequency jumps by the frequency difference between cavity modes. Assuming that the frequency of the mode 1 (or the mode 2) is resonant with (or far-detuned from) the atomic resonance, the optical pumping rate, $R\propto I/(\Delta\nu^{2}+(\Gamma_{L}/2)^{2})$, will exhibit steep changes, as shown in Fig. 1. The change in optical pumping rate is sufficiently sudden, in other words, discontinuous, because the switching time of mode-hopping, on the order of 10 ns, is much shorter than the Larmor period of 2.85 $\mu$s in a magnetic field of $50~{}\mu$T butler2017direct .
In order to analyze the effect of the discontinuous optical pumping induced by a mode-hopping, we investigate the macroscopic spin Bloch equation including the optical pumping modulation term, $R(t)$.
The Bloch equation model, widely used for atomic magnetometers, is suitable for modeling arbitrary modulations distorted by mode-hopping grujic2013atomic . The Bloch equation is given by
@@
Here, $\vec{s}=(e^{*}\times e)/i$ represents the photon spin vector where $e$ represents a unit polarization vector in the yz plane so that $\vec{s}=s\hat{x}$. $\Gamma_{re}$ is the spin relaxation rate, mainly contributed by the spin exchange collisions between 87Rb atoms. $R(t)$ is the optical pumping rate for unpolarized atoms, depending on the laser detuning from the $D_{1}$ resonance and the laser intensity vasilakis2011precision . $\gamma$ is the gyromagnetic ratio. We consider the magnetic field to be directed in the $\hat{y}$ axis.
@@
In the case of a square wave optical pumping modulation, $R_{0}$ varies according to the duty cycle $d$ when the maximum value of $R(t)$ is fixed. Assuming the maximum optical pumping rate of the square wave optical pumping modulation is $R_{m}$, then $S_{0}$ can be expressed by
@@
The maximum $S_{0}$ is 1/2 when $d\rightarrow 0$ in the regime where $\Gamma_{re}\ll R_{m}d$. This implies that a delta-function-like optical pumping induces the largest $S_{0}$ as long as the peak intensity of the laser pulse is strong enough to satisfy $R_{m}\gg\Gamma_{re}/d$. The reason why the delta-function-like optical pumping is optimal can be explained in terms of $\vec{s}$ and $\langle\vec{S}\rangle$. In the Bell-Bloom configuration, optical pumping not only creates the Rb spin polarization along the direction of the pumping beam,but also relaxes the transverse spin. When the direction of $\vec{s}$ does not coincide with $\langle\vec{S}\rangle$, optical pumping destroys the transverse spin precession and aligns the atomic spin with the direction of the pumping beam. It ceases to serve as a transverse spin relaxation mechanism only if $\langle\vec{S}\rangle$ is oriented in the same direction of $\vec{s}$. As a result, we can conclude that the delta-function-like optical pumping is optimal.
In the regime where the optical pumping rate $R_{m}$ is not much larger than $\Gamma_{re}/d$, the nonzero $d$ can be optimal. For example, in the weak pumping condition where $\Gamma_{re}=2\pi\times 685$ Hz, $R_{m}=2\pi\times 300$ Hz, and $|s|=1/2$, the optimal duty cycle is calculated to be around 50 %. In this case, $S_{0}$ of the discontinuous optical pumping with a 50 % duty cycle is $4/\pi$ times larger than that of the sinusoidal optical pumping modulation if the maximum optical pumping rates are the same.
@@ III Experimental Setup
Figure 2: Experimental setup for the single beam all-optical magnetometer. DBR laser: Distributed Bragg Reflector laser, PBS: Polarization Beam Splitter, QWP: Quarter Wave Plate, HWP: Half Wave Plate, BPD: Balanced Photodetector. A burst sine wave modulation to the DBR laser realizes pulsed operation at 1 kHz. The red and blue regions show the pumping and probing stages of the pulsed magnetometer, respectively.
@@
Figure 3: (a) The 87Rb D1 absorption spectra in the atomic vapor cell. The grey dots indicate the absorption of the linearly polarized laser beam, while the black solid line depicts the fitting curve. The Full Width at Half Maximum linewidth $\Gamma_{L}$ is 10.5 GHz which implies that the pressure of $\textrm{N}_{2}$ in our cell is about 580 Torr. The red and blue solid lines depict the absorption spectra of hyperfine transitions from $F_{g}=1$ and $F_{g}=2$, respectively. The black arrows show how the laser frequency detuning changes as the laser current increases. (b) Laser detuning as a function of LD injection currents. The mode-hop current is 67.1 mA. (c) Relative Faraday rotation angle (black solid line) and transverse spin relaxation rate (red solid line) $\Gamma_{re}$ as a function of LD injection current. The injection current below the mode-hop current induces a larger Faraday rotation angle and transverse spin relaxation rate because the corresponding laser frequency is within the 87Rb D1 absorption spectrum. (d) Relative optical power of the laser beam after passing through the atomic vapor cell as a function of LD injection current. The laser injection current above the mode-hop current gives higher beam power because the off-resonant light experiences weak absorption.
@@
The experimental setup is shown in Figure 2. A DBR laser (Photodigm 794.978 nm series with TOSA package) is used as a light source. Mode-hopping in the DBR laser occurs under fixed conditions due to its monolithic nature, making the jump to the output frequency controllable. Two main factors influence the frequency tuning of the DBR laser output: the temperature of the LD heat sink and the laser injection current. When modulating the injection current with a stabilized laser temperature, a slight alteration in laser current triggers a transition between two adjacent cavity modes, as illustrated in Fig. 1. The transition of the lasing mode near the mode-hop current results in a significant jump in the laser output frequency, which was almost 20 GHz in our case. With the aid of mode-hopping, a sinusoidal modulation of the LD injection current is converted to a discontinuous laser output frequency profile without any additional optics.
@@
The linearly polarized laser beam is first transformed into an elliptically polarized beam after passing through a QWP. The optimal QWP angle to increase the signal is 25 and the corresponding $|s|$ is about 0.67 (see the appendix). The laser beam serves both as the probing and pumping beam. It illuminates a $5\times 5\times 5$ m3 cubic cell which contains 87Rb vapor and $\mathrm{N_{2}}$ gas at a pressure of 580 Torr (see Figure 3 (a)).
@@
The homemade atomic gas cell is heated up to 103 C using a two-layer structured pair of AC heaters designed to negate magnetic field noise yim2018note . After passing through the atomic gas cell, which is inside a two layered $\mu$-metal shield, the laser beam goes through a HWP. We note that a four layered $\mu$-metal shield (MS-1, Twinleaf) is applied for sensitivity estimation to further reduce unwanted environment magnetic field noise. A polarimeter, consisting of a PBS and a BPD, measures the degree of polarization rotation of the elliptically polarized beam induced by the paramagnetic Faraday rotation effect. The rotation angle of the polarization is proportional to $\Delta\nu/[\Delta\nu^{2}+(\Gamma_{L}/2)^{2}]$ where $\Delta\nu$ is the laser detuning in the probing stage.
@@
Our pulsed magnetometer alternates between pumping and probing modes at a repetition rate of 1 kHz. The laser injection current is modulated by a burst sine wave generated by a function generator (SDG2082X, Siglent), realizing a pulsed magnetometer operation. In the pumping stage, depicted as the red region in Fig. 2, the LD injection current is modulated at the Larmor frequency of ${}^{87}Rb$ (which is 100 kHz in our case), and incorporates a mode-hop current into the modulation of the LD injection current. The mode-hop current is 67.1 mA when the TEC temperature of the LD is set to 32.05 C. When the LD injection current is above the mode-hop current, the laser output frequency deviates from the Rb resonance line as the laser detuning becomes almost 20 GHz, shown as an arrow in Fig. 3. On the other hand, when the LD current is below the mode-hop current, the laser output frequency gets close to the $F_{g}=1$ transition line as the laser detuning becomes almost 0 GHz. Due to the pressure broadening of 10.5 GHz, the laser beam also weakly induces the transitions from $F_{g}=2$.
@@
In the probing stage, illustrated by the blue region in Fig. 2, the sinusoidal modulation of the LD injection current ceases and the injection current is fixed at a certain value at which the laser frequency is off-resonant. The detuned laser beam undergoes the paramagnetic Faraday rotation effect where the polarization rotation angle is proportional to the atomic spin projected in the direction of the laser beam propagation. The laser beam with rotated polarization is directed into a polarimeter. A damped sine signal on which the FID is imprinted can be obtained from the BPD. It can be described by
@@
In order to estimate the sensitivity of our magnetometer, we apply a frequency counter (53230A, Keysight) which measures the frequency of the signal based on the zero-crossing detection method. The reciprocal of the time between two zero-crossing points becomes the Larmor frequency.
Figure 4: (a) FID amplitude as a function of LD current DC offsets and modulation depths. The results are divided into four regions by two dashed lines (the boundaries where modulation includes a mode-hop current) and a straight line (the mode-hop current at 67.1 mA). (b) A semi-theoretical result. $S_{0}$ is calculated based on the optical pumping profile shown in the top panel. The optical pumping rate profile is extracted by subtracting the minimum relaxation rate from the measured relaxation rate shown in Fig. 3 (c). (c) The semi-theoretical result is based on the step-function-like optical pumping profile. (d) FID signals at four regions. The insets show 1-period temporal pumping rate profiles extracted from the top panel of (b). The modulation depth is fixed at 1.8 mApp.